OpenStudy (anonymous):

@kc_kennylau why do you subtract and then add the sin^2 t ?

sammixboo (sammixboo):

Welcome to Open Study! Having any questions or problems? Just ask me or anyone with a purple A next to their username!

OpenStudy (anonymous):

ok thx! @sammixboo

OpenStudy (kc_kennylau):

What was the question?

OpenStudy (anonymous):

cos^4t - sin^4t = 1-2sin^2t you answered this question asked by @queenbeedoe

OpenStudy (kc_kennylau):

\[\cos^4t-\sin^4t\]\[=(\cos^2t+\sin^2t)(\cos^2t-\sin^2t)\]\[=\cos^2t-\sin^2t\]\[=\cos^2t+\sin^2t-2\sin^2t\]\[=1-2\sin^2t\]

OpenStudy (kc_kennylau):

In order to form \(\cos^2t+\sin^2t\).

OpenStudy (anonymous):

Oh i see but when you were solving for @queenbeedoe you wrote the third step like this: = cos^2t + (sin^2t - sin^2t) - sin^2t so i was confused.

OpenStudy (anonymous):

OHHHHHH i understand it now!!! nevermind im so sorry im so stupid. You wrote it like that to get a "1" for 1- 2sin^2t am i correct?

OpenStudy (kc_kennylau):

Yep :)